# Do You Trust Your Model? Emerging Malware Threats in the Deep Learning Ecosystem

### v2

@@ IV-C Self-Extraction and Execution
The trigger is the mechanism that extracts and executes the malicious payload in the victim’s system. MaleficNet leverages the default serialization/deserialization behavior of popular ML libraries such as PyTorch [15] to automatically extract and execute the embedded malware payload. The default save format for PyTorch models is a zip archive, comprising the serialized model object (which describes the model’s architecture) and the values of each individual tensor (the actual parameters of the model), each saved as a separate entry in the archive. The model architecture description includes the reference to the tensors comprising each layer of the network, which are used while loading the model to retrieve the correct tensor values. The standard deserialization method used to load PyTorch models, including popular models such as the recently released Llama2, relies on the toch.load() function. Internally, this method opens the zip archive, extracts and loads the model architecture description, and progressively retrieves and loads the value of the tensors associated with each parameter in the architectural description. Unfortunately, the deserialization method used to load the model architecture description allows an adversary who has control over the saved model file to inject code that can be executed during the loading phase of the architecture description [16] [^3] PyTorch recently (on January 29, 2025) updated the default parameters of its deserialization function to prevent code execution, after our findings were rendered public. However, self-execution remains possible if non-default parameters are used..
@@ VI Evaluation
TABLE II: Detection rate reported on Metadefender [39] for plain malware binaries, stegomalware version of the malware created using OpenStego [61], and the stegomalware obtained using MaleficNet (Mal. in the table)
@@ VI-A2 Statistical Analysis
We further study whether we can detect MaleficNet models based on the characteristics of the CDMA signal. As discussed in Section IV, MaleficNet employs CDMA to embed a malware payload into the weights of the target model. Since the spreading codes are randomly generated by sampling from $\{-1,1\}$, the CDMA code $C_{j}b_{j}$ follows the binomial distribution (see Section IV). Therefore, we can look for signs of binomiality in the parameters’ distribution as an indicator of manipulation of the weights. We use the two-sample KS test to assess whether the distribution of the MaleficNet model parameters resembles a binomial. For all considered architectures and datasets, our results show that the parameter distribution of the MaleficNet model does not follow a binomial distribution. This result is expected, as the large parameter space of DNNs allows us to embed the payload using only a small fraction of the available parameters. Moreover, the injection is performed with a low-power signal that minimally affects the original model’s parameters. Consequently, the injected signal is not strong enough, nor large enough, to meaningfully alter the original distribution. We summarize our results in Figure 1, which provides a visual comparison between the distribution of the model parameters before and after the malware injection with MaleficNet. As highlighted in the figure, the difference in distribution between the baseline model (without the malware) and the MaleficNet model is minimal.
@@
Malware Bas. Mal. Bas. Mal. Bas. Mal. Bas. Mal. Bas. Mal.
Stuxnet 62.13 61.22 75.69 75.34 76.96 76.87 70.13 70.09 73.37 73.34
Destover 62.13 52.36 75.69 74.89 76.96 76.79 70.13 70.05 73.37 73.28
Asprox - - 75.69 74.76 76.96 76.64 70.13 70.01 73.33 73.22
Bladabindi - - 75.69 74.59 76.96 76.50 69.93 70.04 73.37 73.11
Zeus-Bank - - - - 76.96 76.11 70.13 69.61 73.37 73.02
Eq.Drug - - - - 76.96 75.62 70.13 69.51 73.37 72.89
Zeus-Dec - - - - 76.96 75.24 70.13 69.37 73.37 72.72
Kovter - - - - 76.96 75.01 70.13 69.40 73.37 72.61
Cerber - - - - 76.96 74.51 70.13 69.26 73.37 72.23
Ardamax - - - - - - 70.13 69.12 73.37 72.01
NSIS - - - - - - 70.13 68.99 73.37 71.91
Kelihos - - - - - - 70.13 68.63 73.37 71.72
@@
Malware Bas. Mal. Bas. Mal. Bas. Mal.
Stuxnet 0.41 0.40 0.67 0.65 1.25 1.25
Destover - - - - 1.25 1.25
Asprox - - - - 1.25 1.23
Bladabindi - - - - 1.25 1.23
Zeus-Bank - - - - 1.25 1.21
Eq.Drug - - - - 1.25 1.20
Zeus-Dec - - - - 1.25 1.15
Kovter - - - - 1.25 1.13
Cerber - - - - 1.25 1.10
@@ VI-B1 Traditional Training Setting
Theoretically, using CDMA, we can encode up to one bit per channel use [54] (see Section IV). Nevertheless, this feat is possible when the noise in the channel is Gaussian distributed. In our case, the noise is the learned weights of the neural network which are tightly related to the model’s behavior on the intended task. This restricts the amount of information we can encode in each weight parameter; We should not deteriorate the performance. The amount of information we can encode in a network can not be predicted beforehand due to the fact that the performance of the model relies on the values of the weight parameters learned, and to maintain that performance, those weight parameters can not be altered by a large amount.
@@
Malware Bas. Mal. Bas. Mal. Bas. Mal. Bas. Mal. Bas. Mal.
Stuxnet 98.28 98.05 97.29 97.24 98.15 98.05 98.83 98.78 99.18 98.82
Destover 98.28 97.68 97.29 97.13 98.15 97.91 98.83 98.72 99.18 98.79
Asprox 98.28 97.46 97.29 97.08 98.15 97.68 98.83 98.75 99.18 98.77
Bladabindi - - 97.29 96.74 98.15 97.12 98.83 98.73 99.18 98.76
Zeus-Bank - - - - 98.15 96.18 98.83 97.99 99.18 98.56
Eq.Drug - - - - 98.15 95.98 98.83 97.91 99.18 98.41
Zeus-Dec - - - - 98.15 95.15 98.83 97.79 99.18 98.25
Kovter - - - - 98.15 93.45 98.83 97.85 99.18 98.51
Cerber - - - - - - 98.83 98.22 99.18 98.94
Ardamax - - - - - - 98.83 98.07 99.18 98.42
NSIS - - - - - - 98.83 97.88 99.18 98.32
Kelihos - - - - - - 98.83 96.11 99.18 97.63
@@ VI-B3 Reduced Parameter Width Setting
Malware Bas. Mal. Bas. Mal. Bas. Mal. Bas. Mal. Bas. Mal.
Stuxnet 42.90 42.83 36.40 36.35 51.20 51.18 52.20 52.15 45.30 45.42
Destover 42.90 43.27 36.40 37.44 51.20 52.88 52.20 52.56 45.30 46.27
Asprox 42.90 42.85 36.40 35.45 51.20 51.19 52.20 51.42 45.30 45.06
Bladabindi 42.90 43.34 36.40 36.15 51.20 52.16 52.20 52.62 45.30 45.87
Zeus-Bank 42.90 42.80 36.40 36.12 51.20 51.05 52.20 52.15 45.30 45.33
Eq.Drug 42.90 42.32 36.40 35.88 51.20 50.57 52.20 52.13 45.30 45.01
Zeus-Dec 42.90 42.40 36.40 35.87 51.20 50.78 52.20 52.10 45.30 45.07
Kovter 42.90 41.79 36.40 36.98 51.20 52.52 52.20 52.90 45.30 45.67
Cerber 42.90 43.32 36.40 37.67 51.20 51.97 52.20 52.75 45.30 46.18
Ardamax 42.90 42.68 36.40 37.08 51.20 51.93 52.20 51.45 45.30 45.52
NSIS 42.90 42.50 36.40 35.69 51.20 49.11 52.20 50.37 45.30 44.30
Kelihos 42.90 42.40 36.40 35.45 51.20 50.05 52.20 50.05 45.30 44.35
@@ VI-C Robustness
These findings reinforce that fine-tuning—even across multiple stages—does not significantly deteriorate the MaleficNet payload signal. In each repurposing step, a moderate drop in SNR is observed, typically more noticeable when fine-tuning on complex tasks, which induce larger changes in model parameters to achieve convergence. Despite this, the overall reduction remains limited after each stage. In the analyzed scenario, where the Stuxnet payload was embedded via MaleficNet with an initial SNR slightly above 8:1, the signal remains robust and fully recoverable. Importantly, reducing the SNR to levels where payload bits become unreliable would require a large number of diverse, successive repurposings. This observation aligns with the findings of Torrieri [60], who notes that for successful CDMA extraction, the SNR must remain above 1:1—ensuring that the embedded signal can still be distinguished from background noise and accurately decoded.
@@ VII-C2 Split Learning
The evolving threat landscape for MaleficNet and payload embedding remains a significant concern, highlighting the need for continued research and proactive security measures. Consequently, future work should be dedicated to developing robust defense mechanisms to mitigate attacks such as MaleficNet, ensuring resilience even within the federated learning paradigm. Addressing these challenges will require innovative approaches that enhance security while preserving the decentralized nature of federated systems.
@@ VIII Related Work
Liu et al. [33], to the best of our knowledge, are the first to create a new breed of stegomalware through embedding malware into a deep neural network model. They proposed four methods to embed a malware binary into the model and designed and evaluated triggering mechanisms based on the logits of the model. The first malware embedding method introduced by Liu et al. [33] is LSB substitution, where the malware bits are embedded into the model by replacing the least significant bits of the models’ parameters. The second method consists of a more complex version of the LSB substitution. The idea behind it resides in substituting the bytes of a set of models’ weight parameters with the bytes of the malware payload. After that, they perform model retraining by freezing the modified weight parameters (where the malware is placed) to restore model performance using only the remainder of the weight parameters. Alongside those two methods, Liu et al. [33] proposed mapping-based techniques to map (and sometimes substitute) the values or the sign of the network’s weights to the bit of the malware. They call these methods value-mapping and sign-mapping, respectively. Liu et al. [33] demonstrated that, in the case of LSB substitution, resilience training, and value-mapping, even a single flip in one bit would corrupt the malware, thus rendering these payload embedding techniques unreliable and unusable in practice since even a simple fine-tuning could disrupt the malware extraction. Sign-mapping is the most robust of the four payload embedding techniques proposed in [33], but it suffers from several limitations. Sign-mapping maps the bits of the malware payload to the sign of the model’s weights. This means that the number of bits it can map is more limited than other methods. Based on data reported by Liu et al. [33], the amount of bits that the sign-mapping technique can embed is of the same order of magnitude as the number of bits MaleficNet can embed. Compared to MaleficNet, sign-mapping of Liu et al. [33] requires significantly more information to perform the payload extraction, i.e., the permutation map of the weights. In contrast, MaleficNet extractor only needs to know the seed to generate CDMA spreading codes. Wang et al. [66] proposed fast-substitution as a way to embed malware into a deep neural network. Fast-substitution works by substituting the bits of a selected group of weights in a model with the ones from the malware. In case the performance of the resulted model is highly impaired, the authors restore the performance of the model similar to the resilience training method presented by Liu et al. [33], i.e., freezing the group of weights selected to embed the payload and retrain the model. Fast-substitution suffers the same drawbacks as LSB substitution, resilience training, and value-mapping from Liu et al. [33], making it unusable in the supply-chain attack scenario, where the models are usually fine-tuned. Wang et al. further extended their work [67] proposing two additional techniques: most significant byte (MSB) reservation and half-substitution. Both the methods rely on the fact that the model performance is better maintained when the first bytes of each model weight are preserved. Even if they can guarantee less performance degradation, those methods suffer from the same weaknesses as fast-substitution, making them unsuitable in cases where fine-tuning is used.

### v3

# Paper accepted at IEEE Transactions on Dependable and Secure Computing

@@ IV-C Self-Extraction and Execution
The trigger is the mechanism that extracts and executes the malicious payload in the victim’s system. MaleficNet leverages the default serialization/deserialization behavior of popular ML libraries such as PyTorch [15] to automatically extract and execute the embedded malware payload. The default save format for PyTorch models is a zip archive, comprising the serialized model object (which describes the model’s architecture) and the values of each individual tensor (the actual parameters of the model), each saved as a separate entry in the archive. The model architecture description includes the reference to the tensors comprising each layer of the network, which are used while loading the model to retrieve the correct tensor values. The standard deserialization method used to load PyTorch models, including popular models such as the recently released Llama2, relies on the toch.load() function. Internally, this method opens the zip archive, extracts and loads the model architecture description, and progressively retrieves and loads the value of the tensors associated with each parameter in the architectural description. Unfortunately, the deserialization method used to load the model architecture description allows an adversary who has control over the saved model file to inject code that can be executed during the loading phase of the architecture description [16] [^3] PyTorch recently (on January 29, 2025) updated the default parameters of its deserialization function to prevent code execution, after our findings were rendered public. However, self-execution remains possible if non-default parameters are used. Moreover, to execute the malicious payload hidden within the model does not need to rely only on such vulnerabilities. For example, the attackers might shift their attention to exploiting companion binaries, abusing inter-process communication, or leveraging side-channel vulnerabilities to extract and execute the malware..
@@ VI Evaluation
TABLE II: Detection rate reported on Metadefender [39] for plain malware binaries, stegomalware version of the malware created using OpenStego [61], and the stegomalware obtained using MaleficNet (Mal. in the table).
@@ VI-A2 Statistical Analysis
We further study whether we can detect MaleficNet models based on the characteristics of the CDMA signal. As discussed in Section IV, MaleficNet employs CDMA to embed a malware payload into the weights of the target model. Since the spreading codes are randomly generated by sampling from $\{-1,1\}$, the CDMA code $C_{j}b_{j}$ follows the binomial distribution (see Section IV). Therefore, we can look for signs of binomiality in the parameters’ distribution as an indicator of manipulation of the weights. We use the two-sample KS test to assess whether the distribution of the MaleficNet model parameters resembles a binomial. For all considered architectures and datasets, our results show that the parameter distribution of the MaleficNet model does not follow a binomial distribution. This result is expected, as the large parameter space of DNNs allows us to embed the payload using only a small fraction of the available parameters. We summarize our results in Figure 1, which provides a visual comparison between the distribution of the model parameters before and after the malware injection with MaleficNet. As highlighted in the figure, the difference in distribution between the baseline model and the MaleficNet model is minimal.
Remark: Regarding the stealthiness evaluation, we note that the MaleficNet models remain completely undetected by anti-virus software and the malicious presence in the models weights is unable to be detected even through statistical tests on the parameters distribution.
@@
Baseline 62.13 75.69 76.96 70.13 73.37
Malware Mal. $\Delta$ Mal. $\Delta$ Mal. $\Delta$ Mal. $\Delta$ Mal. $\Delta$
Stuxnet 61.22 -0.91 75.34 -0.35 76.87 -0.09 70.09 -0.04 73.34 -0.03
Destover 52.36 -9.77 74.89 -0.80 76.79 -0.17 70.05 -0.08 73.28 -0.09
Asprox - - 74.76 -0.93 76.64 -0.32 70.01 -0.12 73.22 -0.11
Bladabindi - - 74.59 -1.10 76.50 -0.46 70.04 -0.11 73.11 -0.26
Zeus-Bank - - - - 76.11 -0.85 69.61 -0.52 73.02 -0.35
Eq.Drug - - - - 75.62 -1.34 69.51 -0.62 72.89 -0.48
Zeus-Dec - - - - 75.24 -1.72 69.37 -0.76 72.72 -0.65
Kovter - - - - 75.01 -1.95 69.40 -0.73 72.61 -0.76
Cerber - - - - 74.51 -2.45 69.26 -0.87 72.23 -1.14
Ardamax - - - - - - 69.12 -1.01 72.01 -1.36
NSIS - - - - - - 68.99 -1.14 71.91 -1.46
Kelihos - - - - - - 68.63 -1.50 71.72 -1.65
@@
Baseline 0.41 0.67 1.25
Malware Mal. $\Delta$ Mal. $\Delta$ Mal. $\Delta$
Stuxnet 0.40 -0.01 0.65 -0.02 1.25 0.0
Destover - - - - 1.25 0.0
Asprox - - - - 1.25 -0.02
Bladabindi - - - - 1.25 -0.02
Zeus-Bank - - - - 1.25 -0.04
Eq.Drug - - - - 1.25 -0.05
Zeus-Dec - - - - 1.25 -0.10
Kovter - - - - 1.25 -0.12
Cerber - - - - 1.25 -0.15
@@ VI-B1 Traditional Training Setting
Baseline 98.28 97.29 98.15 98.83 99.18
Malware Mal. $\Delta$ Mal. $\Delta$ Mal. $\Delta$ Mal. $\Delta$ Mal. $\Delta$
Stuxnet 98.05 -0.23 97.24 -0.05 98.05 -0.10 98.78 -0.05 98.82 -0.36
Destover 97.68 -0.60 97.13 -0.16 97.91 -0.24 98.72 -0.11 98.79 -0.39
Asprox 97.46 -0.82 97.08 -0.21 97.68 -0.47 98.75 -0.08 98.77 -0.41
Bladabindi - - 96.74 -0.55 97.12 -1.03 98.73 -0.10 98.76 -0.42
Zeus-Bank - - - - 96.18 -1.97 97.99 -0.84 98.56 -0.62
Eq.Drug - - - - 95.98 -2.17 97.91 -0.92 98.41 -0.77
Zeus-Dec - - - - 95.15 -3.00 97.79 -1.04 98.25 -0.93
Kovter - - - - 93.45 -4.70 97.85 -0.98 98.51 -0.67
Cerber - - - - - - 98.22 -0.61 98.94 -0.24
Ardamax - - - - - - 98.07 -0.76 98.42 -0.76
NSIS - - - - - - 97.88 -0.95 98.32 -0.86
Kelihos - - - - - - 96.11 -2.72 97.63 -1.55
@@ VI-B3 Reduced Parameter Width Setting
Baseline 42.90 36.40 51.20 52.20 45.30
Malware Mal. $\Delta$ Mal. $\Delta$ Mal. $\Delta$ Mal. $\Delta$ Mal. $\Delta$
Stuxnet 42.83 -0.07 36.35 -0.05 51.18 -0.02 52.15 -0.05 45.42 +0.12
Destover 43.27 +0.37 37.44 +1.04 52.88 +1.68 52.56 +0.36 46.27 +0.97
Asprox 42.85 -0.05 35.45 -0.95 51.19 -0.01 51.42 -0.78 45.06 -0.24
Bladabindi 43.34 +0.44 36.15 -0.25 52.16 +0.96 52.62 +0.42 45.87 +0.57
Zeus-Bank 42.80 -0.10 36.12 -0.28 51.05 -0.15 52.15 -0.05 45.33 +0.03
Eq.Drug 42.32 -0.58 35.88 -0.52 50.57 -0.63 52.13 -0.07 45.01 -0.29
Zeus-Dec 42.40 -0.50 35.87 -0.53 50.78 -0.42 52.10 -0.10 45.07 -0.23
Kovter 41.79 -1.11 36.98 +0.58 52.52 +1.32 52.90 +0.70 45.67 +0.37
Cerber 43.32 +0.42 37.67 +1.27 51.97 +0.77 52.75 +0.55 46.18 +0.88
Ardamax 42.68 -0.22 37.08 +0.68 51.93 +0.73 51.45 -0.75 45.52 +0.22
NSIS 42.50 -0.40 35.69 -0.71 49.11 -2.09 50.37 -1.83 44.30 -1.00
Kelihos 42.40 -0.50 35.45 -0.95 50.05 -1.15 50.05 -2.15 44.35 -0.95
@@
Remark: Regarding the performance evaluation, we note that the MaleficNet models exhibit similar performance to their baseline counterparts in different scenarios such as standard, federated learning and even reduced parameter-width setting.
@@ VI-C Robustness
These findings reinforce that fine-tuning—even across multiple stages—does not significantly deteriorate the MaleficNet payload signal. In each repurposing step, a moderate drop in SNR is observed, typically more noticeable when fine-tuning on complex tasks, which induce larger changes in model parameters to achieve convergence. Despite this, the overall reduction remains limited after each stage. In the analyzed scenario, where the Stuxnet payload was embedded via MaleficNet with an initial SNR slightly above 8:1, the signal remains robust and fully recoverable. Importantly, reducing the SNR to levels where payload bits become unreliable would require a large number of diverse, successive repurposings.
Remark: Regarding the robustness evaluation, we note that the malware payload embedded via MaleficNet can not be hampered even when the MaleficNet models undergo different fine-tuning and repurposing stages, always exhibiting an SNR ratio above 1:1—ensuring that the embedded signal can still be distinguished from background noise and accurately decoded.
@@ VII-C2 Split Learning
The evolving threat landscape for MaleficNet and payload embedding remains a significant concern, highlighting the need for continued research and proactive security measures. Consequently, future work should be dedicated to developing robust defense mechanisms to mitigate attacks such as MaleficNet, ensuring resilience even within the federated learning paradigm.
@@ VIII Related Work
Liu et al. [33], to the best of our knowledge, are the first to create a new breed of stegomalware through embedding malware into a deep neural network model. They proposed four methods to embed a malware binary into the model and designed and evaluated triggering mechanisms based on the logits of the model. The first malware embedding method introduced by Liu et al. [33] is LSB substitution, where the malware bits are embedded into the model by replacing the least significant bits of the models’ parameters. The second method consists of a more complex version of the LSB substitution. The idea behind it resides in substituting the bytes of a set of models’ weight parameters with the bytes of the malware payload. After that, they perform model retraining by freezing the modified weight parameters (where the malware is placed) to restore model performance using only the remainder of the weight parameters. Alongside those two methods, Liu et al. [33] proposed mapping-based techniques to map (and sometimes substitute) the values or the sign of the network’s weights to the bit of the malware. They call these methods value-mapping and sign-mapping, respectively.
Liu et al. [33] demonstrated that, in the case of LSB substitution, resilience training, and value-mapping, even a single flip in one bit would corrupt the malware, thus rendering these payload embedding techniques unreliable and unusable in practice since even a simple fine-tuning could disrupt the malware extraction. Sign-mapping is the most robust of the four payload embedding techniques proposed in [33], but it suffers from several limitations. Sign-mapping maps the bits of the malware payload to the sign of the model’s weights. This means that the number of bits it can map is more limited than other methods. Based on data reported by Liu et al. [33], the amount of bits that the sign-mapping technique can embed is of the same order of magnitude as the number of bits MaleficNet can embed. Compared to MaleficNet, sign-mapping of Liu et al. [33] requires significantly more information to perform the payload extraction, i.e., the permutation map of the weights. In contrast, MaleficNet extractor only needs to know the seed to generate CDMA spreading codes.
Wang et al. [66] proposed fast-substitution as a way to embed malware into a deep neural network. Fast-substitution works by substituting the bits of a selected group of weights in a model with the ones from the malware. In case the performance of the resulted model is highly impaired, the authors restore the performance of the model similar to the resilience training method presented by Liu et al. [33], i.e., freezing the group of weights selected to embed the payload and retrain the model. Fast-substitution suffers the same drawbacks as LSB substitution, resilience training, and value-mapping from Liu et al. [33], making it unusable in the supply-chain attack scenario, where the models are usually fine-tuned. Wang et al. further extended their work [67] proposing two additional techniques: most significant byte (MSB) reservation and half-substitution. Both the methods rely on the fact that the model performance is better maintained when the first bytes of each model weight are preserved. Even if they can guarantee less performance degradation, those methods suffer from the same weaknesses as fast-substitution, making them unsuitable in cases where fine-tuning is used.